# Statistical Learning of Distributionally Robust Stochastic Control in Continuous State Spaces

### v1

@@ 3.1 The Current-Action-Aware Case
We begin by clarifying notations and stating the assumptions under which our statistical analysis is carried out. For set $S\subset\mathbb{R}^{d}$, let $\mathrm{diam}(S):=\sup_{x,y\in S}|x-y|$, where ||$\cdot$|| ⋅ | denotes the Euclidean distance. For the CAA adversary case, we assume the following.
@@
    The mapping $(x,a)\rightarrow u^{*}(f(x,a,\cdot))$ uniform $\ell$-Lipschitz, i.e.
    $|u^{*}(f(x,a,y))-u^{*}(f(x^{\prime},a^{\prime},y))|\leq\ell(|x-x^{\prime}|+|a-a^{\prime}|).$
In the following Theorems 1 and 2, $u^{*}$ and $\hat{u}$ are solutions to (2.2) with corresponding Wasserstein and $f_{k}$ adversarial ambiguity sets $\mathcal{P}$ and $\widehat{\mathcal{P}}$ centered at $\mu_{0}$ and $\hat{\mu}$, respectively.
@@
Suppose Assumption 2 is in force. Also, assume that the cost is $c_{\vee}$-bounded; i.e. $\sup_{w,y\in\mathbb{W}}c(w,y)\leq c_{\vee}$. Then, with $\mathcal{P}=\left\{{\mu:W_{c}(\mu,\mu_{0})\leq\delta}\right\}$,
$\left\|\hat{u}-u^{*}\right\|\leq\left(32\sqrt{\pi}\alpha\beta D_{\vee}\sqrt{d_{\mathbb{X}}+d_{\mathbb{A}}+1}+\sqrt{2}\alpha\beta^{3/2}\sqrt{\log\frac{1}{\eta}}\right)n^{-\frac{1}{2}}$
w.p. at least $1-\eta$. Here, $D_{\vee}:=\ell(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\delta^{-1}\beta+1$.
@@
[…]
For notation simplicity, define $k^{\prime}:=k/(k-1)$, $c_{k}(\delta):=(1+k(k1)\delta)^{1/k}$, and $a\vee b=\max\left\{{a,b}\right\}$
@@
$\left\|\hat{u}-u^{*}\right\|\leq 30\beta^{2}n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{D+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+d_{\mathbb{A}})\log n}\right)$
w.p. at least $1-\eta$. Here, $D:=d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3\ell\mathrm{diam}(\mathbb{A})\right)$.
@@ 3.2 The Current-Action-Unaware Case
    The action space $\mathbb{A}$ is finite, equipped with the 0-1 distance; i.e. $d(a,a^{\prime})=\mathds{1}\left\{{a=a^{\prime}}\right\}$.
@@
    The mapping $(x,a)\rightarrow u^{*}(f(x,a,\cdot))$ uniform $\ell$-Lipschitz, i.e.
    $|u^{*}(f(x,a,y))-u^{*}(f(x^{\prime},a^{\prime},y))|\leq\ell(|x-x^{\prime}|+\mathds{1}\left\{{a=a^{\prime}}\right\}).$
Remark.
As we will discuss in Appendix C.4, our proof will generalize to continuum action spaces under additional covering number requirements, yielding $n^{-1/2}$ convergence rates in those settings. However, in this paper, for CAU adversaries, we will focus on the cases where $|\mathbb{A}|<\infty$ to get concrete dependencies on the dimensions, diameters, and the size of the action space. However, it is not clear to us if the same rate can be achieved in the CAU setting if, for example, $\mathbb{A}\subset\mathbb{R}^{d}$ is compact and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$. This is a limitation of this work to be addressed by future research.
In the following Theorems 3 and 4, $\bar{u}^{*}$ and $\hat{u}$ are solutions to (2.3) with corresponding Wasserstein and $f_{k}$ adversarial ambiguity sets $\mathcal{P}$ and $\widehat{\mathcal{P}}$ centered at $\mu_{0}$ and $\hat{\mu}$, respectively.
@@
$\left\|\hat{u}-\bar{u}^{*}\right\|\leq\left(32\sqrt{\pi}\alpha\beta\overline{D}_{\vee}\sqrt{d_{\mathbb{X}}+|\mathbb{A}|+1}+\sqrt{2}\alpha\beta^{3/2}\sqrt{\log{\frac{1}{\eta}}}\right)n^{-1/2}$
w.p. at least $1-\eta$. Here, $\overline{D}_{\vee}:=\ell\mathrm{diam}(\mathbb{X})+2\beta+c_{\vee}\delta^{-1}\beta+1$.
@@
$\left\|\hat{u}-u^{*}\right\|$ $\leq 30\beta^{2}n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\rho)^{2}\left(\frac{c_{k}(\rho)}{c_{k}(\rho)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{\overline{D}+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+|\mathbb{A}|)\log n}\right)$
w.p. at least $1-\eta$. Here, $\overline{D}:=d_{\mathbb{X}}\log\left(1+3\ell\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right)$.
4 Lower Bounds on Finite Sample Minimax Risks
In this section, we study lower bounds on the finite sample minimax risk associated with learning DRSC values for both CAA and CAU adversary models. We demonstrate lower bounds that match the corresponding convergence rate upper bounds specified in the previous section.
@@ 4 Lower Bounds on Finite Sample Minimax Risks
As we will discuss in detail later, the operator $\mathcal{K}$ maps the center $\mu$ of the Wasserstein distance and $f_{k}$-divergence constrained adversarial ambiguity sets to the solution of the Bellman equations (2.2) and (2.3). According to the dynamic programming principles outlined in Appendix A.3, this solution corresponds to the DRSC control value. Therefore, $\mathfrak{M}_{n}(\mathcal{U},\mathcal{K})$ represents the error incurred by the optimal learning algorithm (in terms of uniform performance over all centers of the ambiguity sets in $\mathcal{U}$) for the DRSC control value when the sample size is $n$. To match the rate in the previous upper bounds, it is sufficient to consider $\mathcal{U}$ as the family of Bernoulli distributions with parameter $p\in[0,1]$.
Concretely, to establish our lower bound, we consider an instance with infinitely smooth reward and transition function s.t. the minimax risk associated with learning DRSC values has lower bounds that match the convergence rate of upper bounds in Theorem 1-4.
Sepcifically, let $\mathbb{X}=\mathbb{A}=\mathbb{W}=[-1,1]$, $f(x,a,w)=w$, and $r(x,a)=x$. For fixed controller and adversarial ambiguity set, the value operator is the mapping $\mathcal{K}(\mu)=u^{*}$, where $u^{*}$ is the solution to (2.2) with $\mu_{0}$ replaced by $\mu$. Since $f$ and $r$ are independent of $a$, $u^{*}=\bar{u}^{*}$ is also the solution to (2.3).
@@
Given adversarial ambiguity set $\mathcal{P}$, the solution to (2.2) and (2.3) are
@@
Using this instance, we have the following lower bounds.
Theorem 5 (Lower bound for $W_{2}$-distance-constrained Adversary).
Let $\mathcal{P}=\left\{{\mu:W_{c}(\mu,\mu_{0})\leq\delta}\right\}$ with $c(x,y)=|x-y|^{2}$. The minimax risk of learning $u^{*}$ or $\bar{u}^{*}$ over any $\mathcal{U}\supset\left\{{\mu=p\delta_{\left\{{1}\right\}}+(1-p)\delta_{\left\{{0}\right\}}:p\in[0,1]}\right\}$ is lower bounded by
@@
We state the theorem only in terms of the $W_{2}$ distance. This is just for the convenience of calculations. It is not hard to see from the proof that using a Taylor expansion argument, for $W_{p}$ distances with $c(x,y)=|x-y|^{p}$, we have the same $n^{-1/2}$ rate, matching that in Theorem 1 and 3. Also, upon investigating the proof, one will find that the minimax risk of estimating the value function at one single $x$ has the same lower bound on the rate.
Theorem 6 (Lower bound for $f_{k}$-divergence-constrained Adversary).
Let $\mathcal{P}=\left\{{\mu\ll\mu_{0}:D_{f_{k}}(\mu\|\mu_{0})\leq\delta}\right\}$. Then there exist constants $C_{1}(k,\delta)$ and $C_{2}(k,\delta)$ that only depend on $k,\delta$ s.t. whenever $n\geq C_{1}(k,\delta)$, the minimax risk of learning $u^{*}$ or $\bar{u}^{*}$ over any $\mathcal{U}\supset\left\{{\mu=p\delta_{\left\{{1}\right\}}+(1-p)\delta_{\left\{{0}\right\}}:p\in[0,1]}\right\}$ is lower bounded by

### v2

@@ 4.2 The Current-Action-Aware Case
We begin by clarifying notations and stating the assumptions under which our statistical analysis is carried out. For set $S\subset\mathbb{R}^{d}$, let $\mathrm{diam}(S):=\sup_{x,y\in S}|x-y|$, where $|\cdot|$ denotes the Euclidean distance. For the CAA adversary case, we assume the following.
@@
    The mapping $(x,a)\rightarrow r(x,a,\cdot)+u^{*}(f(x,a,\cdot))$ is uniform $L$-Lipschitz, i.e.
    $|r(x,a,w))-r(x^{\prime},a^{\prime},w))|+|u^{*}(f(x,a,w))-u^{*}(f(x^{\prime},a^{\prime},w))|\leq L(|x-x^{\prime}|+|a-a^{\prime}|).$
We remark that, at an intuitive level, the assumptions of bounded domains and Lipschitz continuity are natural for achieving the $\widetilde{O}(n^{-1/2})$ convergence rate in estimating the robust optimal value. Bounded domains rule out the necessity to learn the behavior of the value function at infinity, which would otherwise require an infinite amount of data. Uniform Lipschitz continuity, on the other hand, reflects the natural smoothness inherent in many physical and economic systems (see [1], [28]). It prevents the value function from exhibiting unbounded oscillations or infinite total variation, phenomena that could significantly deteriorate statistical rates. Together, these conditions ensure that small perturbations in the empirical distribution translate into controlled deviations in the robust value, thereby enabling sharp statistical guarantees.
In the following Theorems 2 and 3, $u^{*}$ and $\hat{u}$ are solutions to (3.2) with corresponding Wasserstein and $f_{k}$ adversarial ambiguity sets $\mathcal{P}$ and $\widehat{\mathcal{P}}$ centered at $\mu_{0}$ and $\hat{\mu}$, respectively.
@@
Suppose Assumption 2 is in force. Also, assume that the cost is $c_{\vee}$-bounded; i.e. $\sup_{w,y\in\mathbb{W}}c(w,y)\leq c_{\vee}$. Then, with $\mathcal{P}=\left\{{\mu:W_{c}(\mu,\mu_{0})\leq\delta}\right\}$, we have that
$\left\|\hat{u}-u^{*}\right\|\leq\left(32\sqrt{\pi}\beta D_{\vee}\sqrt{d_{\mathbb{X}}+d_{\mathbb{A}}+1}+\sqrt{2}\beta^{3/2}\sqrt{\log\left(\frac{1}{\eta}\right)}\right)n^{-\frac{1}{2}}$
w.p. at least $1-\eta$. Here, $D_{\vee}:=L(\mathrm{diam}(\mathbb{X})+\mathrm{diam}(\mathbb{A}))+c_{\vee}\delta^{-1}\beta+1$.
@@
Employing the chaining technique, we eliminate an extra $\log n$ factor. This is at a cost of transforming root-log-diameters (c.f. Theorem 3) into linear diameter dependence. Retaining the $\log n$ allows reducing this to root-$\log$-diameters.
For the case of $f_{k}$-divergence, we can get a similar bound for the estimation error as follows. For notation simplicity, define $k^{\prime}:=k/(k-1)$, $c_{k}(\delta):=(1+k(k-1)\delta)^{1/k},$ and $a\vee b=\max\left\{{a,b}\right\}$.
@@
$\left\|\hat{u}-u^{*}\right\|\leq 30\beta^{2}n^{-\frac{1}{k^{\prime}\vee 2}}c_{k}(\delta)^{2}\left(\frac{c_{k}(\delta)}{c_{k}(\delta)-1}\vee 2\right)\left(\frac{1}{k}+\sqrt{D+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+d_{\mathbb{A}})\log n}\right)$
w.p. at least $1-\eta$. Here, $D:=d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+{d_{\mathbb{A}}}\log\left(1+3L\mathrm{diam}(\mathbb{A})\right)$.
Perhaps surprisingly, Theorems 2 and 3 demonstrate that, even though the goal is to learn the value function uniformly over a continuum of states and actions, and despite the additional adversarial robustness imposed by Wasserstein or $f_{k}$-divergence ambiguity sets, the estimation error $\|\hat{u}-u^{*}\|$ still decays at parametric rates. In particular, they do not suffer from the curse of dimensionality (i.e., rates of the form $n^{-c/|d_{\mathbb{X}}|}$), a phenomenon that typically arises when estimating infinite-dimensional objects. These results implies that, even in infinite state-action spaces and under robustness requirements, policy learning within the CAA setting of our DRSC framework remains statistically efficient.
@@ 4.3 The Current-Action-Unaware Case
    The action space $\mathbb{A}$ is finite, equipped with the 0-1 distance; i.e. $d(a,a^{\prime})=\mathds{1}\left\{{a\neq a^{\prime}}\right\}$.
@@
    The mapping $(x,a)\rightarrow r(x,a,\cdot)+\bar{u}^{*}(f(x,a,\cdot))$ is uniform $L$-Lipschitz
    $|r(x,a,w)-r(x^{\prime},a^{\prime},w)|+|\bar{u}^{*}(f(x,a,w))-\bar{u}^{*}(f(x^{\prime},a^{\prime},w))|\leq L(|x-x^{\prime}|+\mathds{1}\left\{{a\neq a^{\prime}}\right\}).$
As in the CAA setting, boundedness and Lipschitz continuity are natural assumptions that enable efficient uniform learning across the state space and control excessive oscillations in the value function. In the CAU case, however, we additionally assume that the action space $\mathbb{A}$ is finite. We will elaborate on this modeling choice in Remark Remark. Building on Assumption 3, we are now prepared to establish convergence guarantees for the CAU adversary.
In the following Theorems 4 and 5, $\bar{u}^{*}$ and $\hat{u}$ are solutions to (3.3) with corresponding Wasserstein and $f_{k}$ adversarial ambiguity sets $\mathcal{P}$ and $\widehat{\mathcal{P}}$ centered at $\mu_{0}$ and $\hat{\mu}$, respectively.
@@
$\left\|\hat{u}-\bar{u}^{*}\right\|\leq c\left(\beta\overline{D}_{\vee}\sqrt{d_{\mathbb{X}}+|\mathbb{A}|+1}+\beta^{3/2}\sqrt{\log{\frac{1}{\eta}}}\right)n^{-1/2}$
w.p. at least $1-\eta$. Here $c=32\sqrt{\pi}$ and
$\overline{D}_{\vee}:=L\mathrm{diam}(\mathbb{X})+2\beta+c_{\vee}\delta^{-1}\beta+1.$
@@
[…]
$\times\left(\frac{1}{k}+\sqrt{\overline{D}+\log\frac{1}{\eta}+2(d_{\mathbb{X}}+|\mathbb{A}|)\log n}\right)$
w.p. at least $1-\eta$. Here, the parameter
$\overline{D}:=d_{\mathbb{X}}\log\left(1+3L\mathrm{diam}(\mathbb{X})\right)+|\mathbb{A}|\log\left(1+6\beta\right).$
Remark.
As we discuss in Appendix C.4, our proof can be extended to continuum action spaces under additional covering number requirements, yielding $n^{-1/2}$ convergence rates in those settings. In this paper, however, for CAU adversaries, we focus on the case $|\mathbb{A}|<\infty$ to obtain explicit dependencies on the dimensions, diameters, and the size of the action space. It is not clear whether the same parametric rate can be achieved when $\mathbb{A}\subset\mathbb{R}^{d}$ is compact and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$. In such cases, an analysis analogous to the CAA setting yields an upper bound of $O(n^{-1/(2\vee d_{\mathbb{A}})})$, but no matching lower bound is available. The exact rate in this setting remains an open question.
Fortunately, if the controller’s set $\mathcal{Q}$ is finite-dimensional (e.g., parametric, $\mathcal{Q}=\{q_{\theta}\in\mathcal{P}(\mathcal{A}):\theta\in\Theta\subset\mathbb{R}^{d}\}$, as is common when applying policy gradient methods with neural-network-parametrized policies), then even with continuous actions the same parametric convergence rate upper bound continues to hold.
5 Learning Optimal DRSC: Minimax Risks Lower Bounds
In this section, we study lower bounds on the finite sample minimax risk associated with learning the optimal value functions for both CAA and CAU adversary models. We demonstrate lower bounds that match the corresponding convergence rate upper bounds specified in the previous section.
@@ 5 Learning Optimal DRSC: Minimax Risks Lower Bounds
As we will discuss in detail later, the operator $\mathcal{K}$ maps the center $\mu$ of the Wasserstein distance and $f_{k}$-divergence constrained adversarial ambiguity sets to the solution of the Bellman equations (3.2) and (3.3). According to the dynamic programming principles, this solution corresponds to the optimal DRSC value. Therefore, $\mathfrak{M}_{n}(\mathcal{U},\mathcal{K})$ represents the error incurred by the optimal learning algorithm (in terms of uniform performance over all centers of the ambiguity sets in $\mathcal{U}$) for the optimal value when a data set with sample size $n$ is available.
Specifically, we establish a lower bound on $\mathfrak{M}_{n}(\mathcal{U},\mathcal{K})$ for a simple value operator $\mathcal{K}$ and a simple class of probability models $\mathcal{U}$. We construct an instance with infinitely smooth reward and transition functions, where the minimax risk of learning the optimal robust value function achieves lower bounds that match the convergence rate upper bounds in Theorems 1–4. It turns out that to obtain such a matching lower bound, it suffices to let $\mathcal{U}$ be the family of Bernoulli distributions with parameter $p\in[0,1]$.
Concretely, for fixed dynamics $f$, reward $r$, and controller and adversarial ambiguity sets $\mathcal{Q}$ and $\mathcal{P}$, the value operator in the CAA (resp. CAU) case is defined by $\mathcal{K}(\mu)=u^{*}$ (resp. $\mathcal{K}(\mu)=\bar{u}^{*}$), where $u^{*}$ (resp. $\bar{u}^{*}$) is the solution to (3.2) (resp. (3.3)) with $\mu_{0}$ replaced by $\mu$. With this definition, we consider the following DRSC instance.
@@
Let $\mathbb{X}=\mathbb{A}=\mathbb{W}=[-1,1]$, $f(x,a,w)=w$, $r(x,a,w)=x$, and $\mathcal{Q}=\mathcal{P}(\mathcal{A})$. Then the solutions to (3.2) and (3.3) are
@@
Using this instance, we can derive the following lower bounds on the minimax statistical risks in both the Wasserstein and the $f_{k}$ setting.
Theorem 6 (Lower bound under $W_{2}$-distance).
Let $\mathcal{P}=\left\{{\mu:W_{c}(\mu,\mu_{0})\leq\delta}\right\}$ with $c(x,y)=|x-y|^{2}$. The minimax risk of learning $u^{*}$ or $\bar{u}^{*}$ over any
$\mathcal{U}\supset\left\{{\mu=p\delta_{\left\{{1}\right\}}+(1-p)\delta_{\left\{{0}\right\}}:p\in[0,1]}\right\}$
is lower bounded by
@@
We state the theorem only in terms of the $W_{2}$ distance. This is just for the convenience of calculations. It is not hard to see from the proof that using a Taylor expansion argument, for $W_{p}$ distances with $c(x,y)=|x-y|^{p}$, we have the same $n^{-1/2}$ rate, matching that in Theorem 2 and 4. Also, upon investigating the proof, one will find that the minimax risk of estimating the value function at one single $x$ has the same lower bound on the rate.
Theorem 7 (Lower bound under $f_{k}$-divergence).
Let $\mathcal{P}=\left\{{\mu\ll\mu_{0}:D_{f_{k}}(\mu\|\mu_{0})\leq\delta}\right\}$. Then there exist constants $C_{1}(k,\delta)$ and $C_{2}(k,\delta)$ that only depend on $k,\delta$ s.t. whenever $n\geq C_{1}(k,\delta)$, the minimax risk of learning $u^{*}$ or $\bar{u}^{*}$ over any
$\mathcal{U}\supset\left\{{\mu=p\delta_{\left\{{1}\right\}}+(1-p)\delta_{\left\{{0}\right\}}:p\in[0,1]}\right\}$
is lower bounded by